# Simple Hamiltonian dynamics is a powerful quantum processing resource

### v1

@@ VII Participation Ratio (PR)
Figure 9: Plot of the APR vs. $\Delta\mathrm{PR}$ for the ZZ-X model with ten qubits. The blue line and markers are the results of the ZZ-X model, where we set $0\leq\theta_{x}\leq 0.25\pi/$, $\theta_{J}=2\pi$, and $\alpha=1.5$. Here, the red line is the result of the ZZX-Ising model, where we used the same parameters as in Fig. 8. The green line is the result of the XX-Ising model with $g/J_{0}=0.0$ and $J_{0}t=3.5$. Black dash lines indicate the result of $\theta_{x}=0.125\pi$ for ZZ-X model. The red and blue triangle markers show the Haar random and random rotations results. Now (a) and (b) depict the results of the same random initial state sets used in Fig. 8 while (c) shows the results from the MNIST dataset.
@@ VIII Discussion and Conclusion
The analysis in this paper has shown that a Hamiltonian dynamics as the simple as ZZ-interaction without the transversal magnetic field can be used to perform non-trivial classical computational tasks. In particular, unlike the previous quantum QERC reservoir [21], no more than two-partite interactions (either physical or effective) are necessary for the QERC to perform. It is well known that the circuit implementation of Haar measure sampling is costly and hence its circuit depth is long. The Clifford+T model however achieves similar testing accuracy, which is an advantage when the quantum computation is done in the fault-tolerant manner.
For small scale quantum processing, the Ising models can be implemented in many alternate physical platforms with high accuracy. The sampling measurement cost does not follow the cost associated with the statistical reconstruction of the distribution function, instead scaling well with the size of the computational basis. This is an advantage in the implementation.
Finally the ZZ-X model outperformed the Haar sampled unitary operation in both the testing accuracy and the sampling cost at the measurement, which directly indicates that the generation of the complexity and the use of it would be two different things in designing quantum machine learning systems. Our results have shown a new approach to design quantum-classical hybrid machine learning models as well as given an insight into the computational process within quantum reservoirs.
@@ Appendix A Quantum Extreme Reservoir Computing (QERC)
where $\bar{p}=(\sum_{l}^{2^{N}}p_{l})/2^{N}=1/2^{N}$ and $S=\sum_{l}^{2^{N}}(p_{l}-\bar{p})^{2}$ are the mean value and standard deviation, respectively. Next, $\mathbf{u}$ is sent to the classical classifier where we employ the single-perceptron mode (SPM) in which the data transfer is given by $\mathbf{y}=f\left(W\mathbf{x}+\mathbf{b}\right)$. Here $\mathbf{y}$ is the final output of the QERC, $W$ and $\mathbf{b}$ are the weight matrix and bias vector for the SPM. The function $f(\cdot)$ is called an activation function. We employ the softmax activation function [65] in the main text because the QERC is applied to the multi-class classification task. Hence, the final output tells the classification class of the input data.
@@
Figure 10: Schematic diagram of the computational process for Quantum Extreme Reservoir Computation (QERC). The bright-blue and bright-green areas represent the quantum and classical processors, respectively. Classical information is encoded into a quantum system using single qubit rotation gates with angles $\theta_{l},\phi_{l}$, where $l=(1,2,\cdots,N)$ and $N$ is the total number of qubits of the quantum processor. In QERC, one can use quantum many-body dynamics (QMD), such as the Ising models discussed in the main text, or a quantum gates sequence as a reservoir. Probability distributions are then determined at the measurement layer and sent to a single-layer classical neural network. The training process is only applied to the classical neural network, while the quantum reservoir is fixed.
@@ B.1 Estimation of Sampling Number
Let us consider a quantum system with $N$ qubits and run it totally $N_{\text{s}}$ times to estimate the probability distribution. We measure the system with a computational basis set. In other words, for each running, we can get a bit string $x\in\mathbb{S}_{N}$, where the set $\mathbb{S}_{N}$ is a set of whole possible N-bits string: $\mathbb{S}_{N}=\left\{00\cdots 0,\cdots,11\cdots 1\right\}$. To simply the discussion, let us assume that we measure a bit string $x$ with $0\leq N_{x}\leq N_{s}$ times. In this assumption, the probability that we get the result is given by
@@
$\langle\tilde{p}_{x}\rangle=\left\langle\frac{N_{x}}{N_{\text{s}}}\right\rangle=p_{x}$ (20)
its indicates that $\tilde{p}_{x}$ is the unbiased estimator, and its variance is given by
@@
As you can see, the error $\epsilon_{\tilde{p}_{x}}$ is dependent on $p_{x}$. As an example, let us consider the two extreme cases. First, let us consider that the probability of the single-bit string $x$ is precisely one. In this case, its error is $\epsilon_{\text{L}}=0$. Next, consider a case where the probability distribution is uniform: $p_{x}={1}/{2^{N}}$ for arbitral bit string $x$. Its error can be $\sim 1/\sqrt{2^{N}N_{s}}$ for the large $N$. We see that it becomes small as the number sampling $N_{s}$ increases. As a state is localized, its estimation error gets smaller. While a state is ergodic, its error gets larger.
Figure 11: Plots show the sampling cost of QERC with different quantum reservoirs: (a) show the sampling number dependence on Kullback–Leibler divergence (KLD) for different models with $N=10$ and $N=12$. (b)-(e) shows sampling number dependence on the performance with different reservoirs (ZZ-Ising, XX-Ising, the Haar random and ZZ-X model), where we set $g/J_{0}=1.0$ for the ZZ-Ising, $g/J_{0}=0.0$. The color lines show the accuracy rate for the probability distribution by assuming one can get the ideal one. (b) to (e) show the sampling number dependence on the performance for the different reservoirs with $N=10$ and $N=12$.
@@
In general, finding the proper value of the constant $C$ is difficult, but we can get an estimate of it as follows. After the measurement, the probability distribution is renormalized to have an average zero to the variance one, as follows,
@@
Now, it is natural to assume that when the above error is much smaller than that of the distribution of $q_{x}$ generated by data, QERC works appropriately. Thus, we assume that the stochastic fluctuation is much smaller than the valiance of the normalized distribution, which is one. It can be written as $D\cdot\mathrm{PR}\cdot\epsilon_{{p}_{x}}<1$, where $D\gg 1$ is a constant that relates to a sensitivity of the classical neural network. Finally, consider a case where $\epsilon_{p_{x}}\sim 1/\sqrt{2^{N}N_{s}}$ is the worst case, and the number of shots can be given ${D^{2}}\cdot\mathrm{PR}^{2}/{2^{N}}<N_{s}$. It indicates that the sampling cost is less as the PR is smaller.
@@ B.2 Numerical results with different systems
First, to see the reconstruction performance of the sampling distribution, we calculate the Kullback-Leibler divergence (KLD) between a theoretical distribution and an empirical distribution, which is numerically obtained by random sampling. The results of KLD with different sizes are shown in Fig. 11 (a) .It means that as system size increases, one needs more sampling to achieve the same level as is well known.
Next, Figs. 11 (b) to (e) show the difference of the required number of sampling for each quantum reservoir with respect to the number of qubits. There would be a trade off between the advantages (the larger Hilbert space and the larger number of the PCA components) and the disadvantage (the statistical overhead of sampling) for the larger number of qubits. The quantum reservoir with a larger number of qubits would perform better at least at up a certain size in theory due to the larger number of the PCA components can be input and the calculation space is larger, however, due to the reconstruction of the amplitude distribution, it may require a larger number of sampling. The latter can be seen in the small number sampling regime in Figs. 11 (c) and (d). This indicates that the difference between the amplitude distribution to be distinguished is more detailed with these quantum reservoirs. By contrast, the ZZ-X model does not have the region where the smaller quantum reservoir achieves a higher accuracy. This is evidence that the sampling in QERC does not necessarily follow statistical scaling, which is a significant advantage for the implementation.

### v2

@@ VII Participation Ratio (PR)
Figure 9: Plot of the APR vs. $\Delta\mathrm{PR}$ for the ZZ-X model with ten qubits. The blue line and markers are the results of the ZZ-X model, where we set $0\leq\theta_{x}\leq 0.25\pi/$, $\theta_{J}=2\pi$, and $\alpha=1.5$. Here, the red line is the result of the ZZX-Ising model, where we used the same parameters as in Fig. 7. The green line results from the XX-Ising model with $g/J_{0}=0.0$ and $J_{0}t=3.5$. Black dashed lines indicate the result of $\theta_{x}=0.125\pi$ for the ZZ-X model. The red and blue triangle markers show the Haar random and random rotations results. Now (a) and (b) depict the results of the same random initial state sets used in Fig. 7 while (c) shows the results from the MNIST dataset.
@@ VIII Discussion and Conclusion
The analysis in this paper has shown that Hamiltonian dynamics as simple as ZZ-interaction without the transversal magnetic field can be used to perform non-trivial classical computational tasks. In particular, unlike the previous quantum QERC reservoir [21], no more than two-partite interactions (either physical or effective) are necessary for the QERC to perform. It is well known that the circuit implementation of Haar measure sampling is costly, and hence its circuit depth is long. The Clifford+T model, however, achieves similar testing accuracy, which is an advantage when the quantum computation is done in a fault-tolerant manner.
For small scale quantum processing, the Ising models can be implemented in many alternate physical platforms with high accuracy. More importantly, it circumvents the existing criticism for the exponential cost of the sampling measurement due to the fact that the sampling cost does not follow the cost associated with the statistical reconstruction of the distribution function. In fact, the accuracy rate converges around the fixed sampling number for $N=10$ and $12$. This is an advantage in the implementation. Additionally, we have confirmed that in other image datasets, such as Fashion-MNIST [75] and K-MNIST [76], we obtained similar results discussed in this paper.
Finally the ZZ-X model outperformed the Haar sampled unitary operation in both the testing accuracy and the sampling cost at the measurement, which directly indicates that the generation of the complexity and the use of it would be two different things in designing quantum machine learning models. Our PR analysis shows that the two main processes in quantum algorithms, information scrambling and the convergence to a solution, can be realized by a simple quantum reservoir such as the ZZ-X model without any programming and parameter optimizations. Our results have shown a new approach to design quantum-classical hybrid machine learning models as well as given an insight into the computational process within quantum reservoirs.
@@ Appendix A Quantum Extreme Reservoir Computing (QERC)
where $\bar{p}=(\sum_{l}^{2^{N}}p_{l})/2^{N}=1/2^{N}$ and $S=\sum_{l}^{2^{N}}(p_{l}-\bar{p})^{2}$ are the mean value and standard deviation, respectively. We refer to the measurement and the standardization together as the M-layer. Next, $\mathbf{u}$ is sent to the classical classifier where we employ the linear classifier in which the data transfer is given by $\mathbf{y}=f\left(W\mathbf{x}+\mathbf{b}\right)$. Here $\mathbf{y}$ is the final output of the QERC, $W$ and $\mathbf{b}$ are the weight matrix and bias vector. The function $f(\cdot)$ is called an activation function. We employ the softmax activation function [77] in the main text because the QERC is applied to the multi-class classification task. Hence, the final output tells the classification class of the input data.
@@
Figure 10: Schematic diagram of the computational process for Quantum Extreme Reservoir Computation (QERC). The bright-blue and bright-green areas represent the quantum and classical processors, respectively. Classical information is encoded into a quantum system using single qubit rotation gates with angles $\theta_{l},\phi_{l}$, where $l=(1,2,\cdots,N)$ and $N$ is the total number of qubits of the quantum processor. In QERC, one can use quantum many-body dynamics (QMD), such as the Ising models discussed in the main text, or a quantum gates sequence as a reservoir. Probability distributions are then determined at the measurement, and data processing is performed (M-layer). Then, the distribution is sent to a linear classifier with the softmax function. The training process is only applied to the classical neural network, while the quantum reservoir is fixed.
@@ B.1 Estimation of Sampling Number
Let us consider a quantum system with $N$ qubits and run it totally $N_{\text{s}}$ times to estimate the probability distribution. We measure the system with a computational basis set. In other words, for each run, we can get a bit string $x\in\mathbb{S}_{N}$, where the set $\mathbb{S}_{N}$ is a set of all possible N-bit strings: $\mathbb{S}_{N}=\left\{00\cdots 0,\cdots,11\cdots 1\right\}$. To simplify the discussion, let us assume that we measure a bit string $x$ with $0\leq N_{x}\leq N_{s}$ times. In this assumption, the probability that we get the result is given by
@@
$\langle\tilde{p}_{x}\rangle=\left\langle\frac{N_{x}}{N_{\text{s}}}\right\rangle=p_{x}.$ (21)
It indicates that $\tilde{p}_{x}$ is the unbiased estimator, and its variance is given by
@@
As you can see, the error $\epsilon_{\tilde{p}_{x}}$ is dependent on $p_{x}$. As an example, let us consider the two extreme cases. First, let us consider that the probability of the single-bit string $x$ is precisely one. In this case, the error is $\epsilon_{\text{L}}=0$ Next, consider a case where the probability distribution is uniform: $p_{x}={1}/{2^{N}}$ for arbitral bit string $x$. Its error can be $\sim 1/\sqrt{2^{N}N_{s}}$ for the large $N$. It becomes small as the number of samples $N_{s}$ increases. As a state is localized, its estimation error gets smaller. While a state is ergodic, its error gets larger.
Figure 11: Plots show the sampling cost of QERC with different quantum reservoirs: (a) shows the sampling number dependence on Kullback–Leibler divergence (KLD) for different models with $N=10$ and $N=12$. (b)-(e) shows sampling number dependence on the performance with different reservoirs (ZZ-Ising, XX-Ising, the Haar random and ZZ-X model), where we set $g/J_{0}=1.0$ for the ZZ-Ising, $g/J_{0}=0.0$. The color lines show the accuracy rate for the probability distribution by assuming one can get the ideal one. (b) to (e) show the sampling number dependence on the performance for the different reservoirs with $N=10$ and $N=12$.
@@
In general, finding the proper value of the constant $C$ is difficult, but we can estimate it as follows. After the measurement, the probability distribution is renormalized to have an average zero to a variance one, as follows,
@@
Now, it is natural to assume that when the above error is much smaller than that of the distribution of $q_{x}$ generated by data, QERC works appropriately. Thus, we assume that the stochastic fluctuation is much smaller than the variance of the normalized distribution, which is one. It can be written as $D\cdot\mathrm{PR}\cdot\epsilon_{{p}_{x}}<1$, where $D\gg 1$ is a constant that relates to the sensitivity of the classical neural network. Finally, consider a case where $\epsilon_{p_{x}}\sim 1/\sqrt{2^{N}N_{s}}$ is the worst case, and the number of shots can be given ${D^{2}}\cdot\mathrm{PR}^{2}/{2^{N}}<N_{s}$. This indicates that the sampling cost is lower as the PR is smaller.
@@
Figure 12: To investigate the number of samples required by QERC scales, we evaluated the absolute difference between the test performance using the theoretical probability distribution and the test performance using the empirical distribution (constructed from shots). (1) shows the absolute difference between the rescaled axis of the number of samples and the Hilbert space dimension, giving $\sqrt{2^{N}/N_{s}}$, where $N_{s}$ is the number of samples. (2) shows the case where the rescaled axis is $\sqrt{N^{2}/N_{s}}$.
@@ B.2 Numerical results with different systems
First, to see the reconstruction performance of the sampling distribution, we calculate the Kullback-Leibler divergence (KLD) between a theoretical distribution and an empirical distribution, which is numerically obtained by random sampling. The results of KLD with different sizes are shown in Fig. 11 (a) .It means that as system size increases, one needs more sampling to achieve the same level, as is well known.
Next, Figs. 11 (b) to (e) show the difference in the required number of samples for each quantum reservoir with respect to the number of qubits. There would be a trade-off between the advantages (the larger Hilbert space and the larger number of PCA components) and the disadvantage (the statistical overhead of sampling) for the larger number of qubits. The quantum reservoir with a larger number of qubits would perform better at least at up a certain size in theory due to the larger number of the PCA components can be input and the calculation space is larger, however, due to the reconstruction of the amplitude distribution, it may require a larger number of sampling. The latter can be seen in the small number sampling regime in Figs. 11 (c) and (d). This indicates that the difference between the amplitude distributions to be distinguished is more detailed with these quantum reservoirs. By contrast, the ZZ-X model does not have the region where the smaller quantum reservoir achieves a higher accuracy. This is evidence that the sampling in QERC does not necessarily follow statistical scaling, which is a significant advantage for the implementation.
To better understand the sample size needed in QERC, we investigated two types of reservoirs: a single ZZX circuit (simply referred to as ZZX) and a model having ten ZZX circuits (denoted as ZZX10). The number of qubits varied from 8 to 16. The first circuit corresponds to the ZZX model described in the main text, where the probability distribution is localized in contrast to other models like the Haar reservoir. The second circuit, ZZX10, applies the ZZX circuit ten times and structurally resembles the discrete-time crystal model introduced in Ref. [21]. Given that ZZX10 alternates between interactions and rotations, its output probability distribution is expected to be adequately spread out. Indeed, for all qubit numbers, the participation ratio (PR) and Shannon entropy of the probability distribution generated by the ZZX10 circuit closely match those of the probability distribution produced by the Haar reservoir.
The numerical results are shown in Fig. 12. The top panel shows the case where the scaling is $\sqrt{{2^{N}/N_{s}}}$. It is clear that the lines corresponding to different qubit numbers do not follow the same curve. This indicates that the required number of qubits does not increase exponentially. While Fig. 12 (2) suggests that the necessary number of shots is more likely to be related to $N^{2}$.
A crucial point is that this scaling law holds regardless of whether the probability distribution is localized or expanded. Additionally, as pointed out in the main text, the ZZX model exhibits a more localized probability distribution than other reservoirs, implying that fewer samples are required than in other models. The common trend for ZZX shows a noticeably smaller slope than that for ZZX10, indicating that the ZZX model requires fewer samples than ZZX10 to achieve similar classification accuracy.